Project

General

Profile

MICE Update 17 May 2015

Progress update from Ken Long
Added by Kyberd, Paul almost 2 years ago

Dear All,

Progress update:

Step IV:

The material for the north PRY arrived at RAL on Tuesday 12May15. When work to install nine (of ten) cryocoolers and the power and sensor cables to the spectrometer solenoids had been completed on Wednesday 13May15, the north PRY legs were installed on Thursday 14May15. The legs were aligned and bolted down on Friday.

In parallel, work on the absorber/focus-coil module continued. The module is now almost ready to be placed on the beam line. The bellows that link the focus-coil module to the spectrometer solenoids have been prepared and will be installed at the start of next week. The focus-coil module will then be moved onto the beam line.

The target team have spliced and terminated the new fibres at junction boxes inside and outside the synchrotron. Work over the coming days will ensure the integrity of the new fibres so that the target may be operated with them.

Progress on LH2:

Excellent progress has been made on preparing the LH2 system for Step IV operation. The integrity of the safety windows has been proven to the satisfaction of the MICE/ISIS H2 working group. Further, the H2 exhaust path has been agreed. These are essential (necessary but not sufficient) steps towards the operation of Step IV with LH2.

The exhaust path now has to be proved and the system commissioned with He gas. In addition, the necessary documentation must be finalised to allow the MICE/ISIS Safety Committee to recommend to the Directors of ISIS and PPD that the LH2 system is fit for operation.

At present we have decided that the first absorber to be studied at Step IV will be LiH. Only once the technical, safety and formal sign-off milestones outlined above are completed can this decision be revisited.

The near future:

MICE Hall:

The big task next week is the installation of the north PRY and the trial assembly of the PRY end plates. The line will be made ready on Monday and Tuesday. Over this period the bellows will be fitted, the absorber/focus-coil module will be placed on the beam line and the channel will be bolted together. On Wednesday, the installation of the north PRY will start. It will be important to complete the installation of the PRY next week to avoid a stack-up of work towards the end of the month.

The last of the cryocoolers in the downstream solenoid will be fitted in the week of the 25th May2015. At this time a short to ground that has been observed in the current leads of the downstream solenoid will also be rectified.

Operations:

The operations plan prepared by the Physics group for ISIS cycle 2015/01a (June) was accepted by the EB at its meeting on 08May15. The Operations group is now working on implementing the plan.

Operations in June and July will be devoted to bringing the spectrometer solenoids and focus-coil module up to operating current, commissioning the tracker readout with beam and performing initial alignment studies of the Step IV apparatus. In June there will be one data-taking shift per day. The shift will be staffed by two shifters and run from 22:00 to 06:00. A “floating shifter” will be in place for the day time (06:00 to 22:00) to support experts in commissioning. Over this period, priority will be given to magnet training and tracker-readout commissioning.

From cycle 2015/01b (July), 24/7 operations (three shifts of two shifters each) will begin. While still scheduled as a commissioning period, it will be the ambition to move more and more to a routine data-taking mode.

In the light of all this excellent progress, I think we can agree that Step IV operations can formally be scheduled to begin in ISIS cycle 2015/02 (August 2015).

Announcements:

1. The MICE instance of the ISIS User Database is now in production. The link to it is:
https://users.facilities.rl.ac.uk/auth/Login.aspx?ReturnUrl=%2fvisits%2f
⁃ Please take a moment to register or to make sure that your existing registration is up to date.
⁃ This is important as information such as your training status is held on the database. Keeping this up to date is part of our safety case and will be scrutinised by ISIS as we seek permission to operate Step IV.
2. The WWW site of CM42 is now live. You can reach it by following the link from the MICE home page:
http://mice.iit.edu
⁃ Please register!

⁃ With best wishes as always …

⁃ Ken


Comments